OpenStudy (anonymous):

given the points (3,-4) and (6,-3) where ax+by=5 find the value of a and b

OpenStudy (anonymous):

Answer: a=5/3 ;b=0

OpenStudy (anonymous):

Kihaga can u show the calculations please

OpenStudy (anonymous):

click on "give medal" for the link to the solution

OpenStudy (anonymous):

Let 3a-4b=5............(i) 6a-3b=5............(ii) solve simultenous eqn by substution method 3a=5+4b...........(iii) from eqn(ii) 10+8b-3b=0 b=-2 3a-4(-2)=5 a=-1 ooh! am sorry i waz too fast! am impassive learner.

OpenStudy (anonymous):

Hahaha that is funy deeprony7. Gonna give you a medal :)